OpenStudy (anonymous):

Find the product of (x + 9i)^2?

OpenStudy (anonymous):

just apply the (a+b)^2 formula (a+b)^2=a^2+b^2+2ab then you will get it....

OpenStudy (anonymous):

do i just plug it into that formula? @chetan552

OpenStudy (solomonzelman):

yes, x for a and 9i for b

OpenStudy (anonymous):

ok thank you @SolomonZelman

OpenStudy (anonymous):

Yes....@mariacruz07

OpenStudy (anonymous):

Sorry, i forgot to say after appling the formula you will see i^2 replace i^2 with -1 becouse "i" means imaginary number... i^2=-1

OpenStudy (anonymous):

@chetan552 would i apply the same formula to find the product of (x-2i)^2?

OpenStudy (phi):

people use the FOIL method to remember how to multiply two binomials (a+b)(c+d) First: a*c Outer a*d Inner: b*c Last: b*d now add up to get ac+ad+bc+bd if you have (a+b)(c -d) treat it like (a+b)(c+ -d) and do the same thing First: a*c Outer: a*-d Inner: b*c Last: b* -d put it all together: ac -ad +bc -bd

OpenStudy (anonymous):

Yes you can....@mariacruz07